OpenStudy (anonymous):

find the zeros of the polynomial function and state the multiplicity of each. f(x)=4(x+7)^2(x-7)^3

OpenStudy (jdoe0001):

\(\bf f(x)=4(x+7)^2(x-7)^3\qquad \textit{set the function to 0}\\ \quad \\ 0=4(x+7)^2(x-7)^3\implies 0=4{\color{blue}{ (x+7)(x+7)}}{\color{red}{ (x-7)(x-7)(x-7)}}\\ \quad \\ \textit{solve for "x" to get the zeros}\quad multiplicity\quad {\color{blue}{ 2}}\ and \ {\color{red}{ 3}}\)

OpenStudy (anonymous):

how is it 2 and 3?

OpenStudy (jdoe0001):

multiplicity = repeats = repetition

OpenStudy (anonymous):

I do not believe it is two and three

OpenStudy (anonymous):

What number would you multiply by another number to get zero?

OpenStudy (anonymous):

anything times zero?

OpenStudy (anonymous):

Correct. So we set the function to 0. And we have 0 = 4(x+7)^2 x (x-7)^2 can you look at it and tell how you are going to be able to create a zero in the equation?

OpenStudy (anonymous):

x=0?

OpenStudy (anonymous):

If x = 0 then it would be 4(0+7)^2 X (0-7) ^ 2 = 9604 not 0. How do you make (x+7) and (x-7) = 0?

OpenStudy (anonymous):

x cannot equal 7 or -7

OpenStudy (anonymous):

You got it. But they want you to "find the zeros of the polynomial" So the answer is 7 and -7...Great job

OpenStudy (anonymous):

thank you!